OpenStudy (anonymous):

multiply (x+1)(4x+3)...does it equal 4x^2+7x+3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

It does indeed. 4x*x = 4x^2 x*3 = 3x 1*4x = 4x 1*3 = 3 Final Answer: 4x^2 + 7x + 3 <== Exactly what you have. Good job! :)

OpenStudy (anonymous):

thx ppl!!

OpenStudy (anonymous):

np